OpenStudy (anonymous):

The length of a rectangle in inches is sixteen less than seven times the width. If the perimeter of the rectangle is 48 inches, what is the length, in inches? A.5 B.8 C.19 D.40

OpenStudy (anonymous):

Let the width be x Perimeter = length + width + length + width = 7x-16 + x + 7x-16 + x = 16x-32 16x-32 = 48 16x = 48+32 = 80 x = 5 Length = 7x-16 = 35-16 = 19 Ans: C

OpenStudy (anonymous):

Thank you. (:

OpenStudy (anonymous):

No problem :)